The local nonprofit youth symphony is planning a concert fundraiser. The organization estimates that 600 tickets can be sold for $15 per person. The fixed costs are $708. The local chamber of commerce office will process ticket orders for a fee of $3 per ticket, to relieve the youth symphony of this responsibility.

(1) How many tickets does your organization have to sell to break even?
(2) How many tickets does your organization have to sell to earn a profit of $4,800?

(3) How much must your organization have in sales dollars to break even (rounded to the nearest cent)?
(4) How much must your organization have in sales dollars to earn a profit of $4,800 (rounded to the nearest cent)?

(5) What is the organization's margin of safety in units and in sales dollars?

SHOW YOUR WORK
Answer #1

Solution 1:

Contribution margin per ticket = $15 - $3 = $12 per ticket

Fixed costs = $708

Nos of ticket to be sold to break even = $708 / $12 = 59 tickets

Solution 2:

Nos of tickets to be sold to earn target profit = (Fixed costs + Target profit) / CM per unit = ($708 + $4,800) / $12 = 459 tickets

Solution 3:

Break even sales in dollar = 59*$15 = $885

Solution 4:

Sales in dollar to earn target profit of $4,800 = 459*$15 = $6,885

Solution 5:

Current sales units = 600

Margin of safety in units = Current sales units - Break even sales units = 600 - 59 = 541 ticket

Margin of safety in dollars = 541*$15 = $8,115